# Artificial Data Point Generation in Clustered Latent Space for Small Medical Datasets

### v1

@@ Abstract
One of the growing trends in machine learning is the use of data generation techniques, since the performance of machine learning models is dependent on the quantity of the training dataset. However, in many medical applications, collecting large datasets is challenging due to resource constraints, which leads to overfitting and poor generalization. This paper introduces a novel method, Artificial Data Point Generation in Clustered Latent Space (AGCL), designed to enhance classification performance on small medical datasets through synthetic data generation. The AGCL framework involves feature extraction, K-means clustering, cluster evaluation based on a class separation metric, and the generation of synthetic data points from clusters with distinct class representations. This method was applied to Parkinson’s disease screening, utilizing facial expression data, and evaluated across multiple machine learning classifiers. Experimental results demonstrate that AGCL significantly improves classification accuracy compared to baseline, GN and kNNMTD. AGCL achieved the highest overall test accuracy of 83.33% and cross-validation accuracy of 90.90% in majority voting over different emotions, confirming its effectiveness in augmenting small datasets.
@@
Synthetic Data Generation, Data augmentation, Small datasets, Parkinson’s Disease Screening, K-means clustering, Normal Distribution
@@ I Introduction
The performance of machine learning models heavily relies on the training dataset [1]. However, collecting a large amount of data is time-consuming and demands considerable resources. Unfortunately, many medical applications do not have enough data to properly train machine learning models. In many practical scenarios, where annotating labels extensively is labor-intensive or only limited datasets are accessible, machine learning models are prone to overfitting and poor generalization [2, 3].
In strict small dataset scenarios, the learner has access to a small number of labeled examples from each class, and the number of classes can also be quite small. In the fields of transfer learning, few-shot learning, and semi-supervised learning, several methods have been proposed. However, learning using small datasets is different from all three research areas. Methods designed to address strict small samples scenarios cannot rely on transfer learning from large amounts of peripheral data. The problem can be notably alleviated by imposing a strong prior on the model. Unfortunately, in many medical applications, we face an unknown domain where such prior knowledge is not available. In the few-shot scenario, the learner has access to a large number of labeled examples from classes not participating in the current classification task, while in the semi-supervised scenario, the learner typically has access to a large number of unlabeled examples [4, 5, 6].
Data augmentation and data generation techniques are effective strategies to address the issues of data scarcity [7]. Generating new synthetic samples is a valuable approach that enables models to generalize better and achieve higher accuracy when encountering unseen data. Leveraging generative adversarial networks to enhance a limited training dataset is particularly attractive, especially now that highly powerful deep generative models are increasingly available. However, these models typically demand a substantial sample size to train effectively, leading to poor performance in scenarios with limited data [8].
In this paper, we introduce the Artificial Data Point Generation in Clustered Latent Space (AGCL) as a solution to address the challenging problems of small medical datasets. The proposed method can be implemented as follows: first, perform feature extraction and then apply K-means clustering. We then assess the results of the clustering for class separation and, if necessary, perform re-clustering. Next, to introduce variety into the dataset, we create additional synthetic data points using the normal distribution based on each cluster’s parameters. Finally, the updated dataset, which combines original and synthetic data, is used to train machine learning models for improved classification performance.
The aim of this study is to tackle the issue of limited training data in the medical field, which makes it challenging to develop an effective classification model. While our primary focus in this paper is on Parkinson’s screening, our methodology can easily be applied to augmenting data in various other fields.
The rest of this article is structured as follows. Section II reviews related work. Section III outlines the methodology. Section IV describes the Dataset which we used for our experiments. Section V presents simulation results and discussions. Finally, Section VI concludes with key takeaways.
Figure 1: Illustrating the AGCL framework
@@ II Related Work
When it comes to working with small datasets, data augmentation appears to be one of the few unmatched methods to boost the efficiency of machine learning models. There are many augmentation methods used for image classification, including random rotation, mirroring, and the addition of Gaussian noise [9]. Heuristic methods such as Cutout [10] partially occlude a square area in each training process to affect the learned features, and Random Erasing [11] partially covers or replaces certain areas of an image. Although these methods are very useful, they may not yield the best results on very small datasets and can negatively affect classification. Blending is another significant method of data augmentation that involves combining images or regions of interest within a single image. For example, the CutMix [12] technique replaces deleted pixels with portions of other images to create a good combination. However, problems can arise with methods such as CutMix, as the pasted patches can clash with the background of the original image and differ semantically from it, causing confusion during label assignment in the training process.
Automatic augmentation [13] has seen significant advancements, focusing on enhancing augmentation techniques through the use of reinforcement learning. This approach aims to automate the selection and fine-tuning of augmentation strategies, promising efficiency gains. However, a key challenge lies in the computational demands of the search algorithm to identify the optimal augmentation approach, particularly as data dimensions and model parameters increase. Feature augmentation represents another common approach to data enhancement, focusing on expanding learned features rather than the input space of the data. FitMatch [14] is one of the methods in this category which uses learned features to create different and elaborate transformations. Since FitMatch increases the number of features, this overloads the data dimensionality and the complexity of the models that would take more time to train when costs are incurred.
Addressing the problem of small datasets, the Least-Square Generative Adversarial Network (LS-GAN) was proposed. This algorithm begins with the independence test on feature removal by correlation measures and eliminates features that are linearly associated with other features. Next the feature selection method of Burato is used which employ random forest algorithm for the enhancement of feature selection. Yet another and rather important aspect of the generative network in the adversarial least squares sense is that the least squares cost function is utilized in this case, which do not allow vanishing gradient to come through [15].
Sivakumar et al. [16] have introduced a modified mega-trend diffusion to address the issues of generating artificial samples for small datasets. The algorithm involves identifying the k nearest neighbors for each sample in the dataset using the k-Nearest Neighbor and then applying Mega-Trend Diffusion to estimate the domain ranges of the neighboring samples. This is a multifunctional approach that can be used to provide artificial data needed in other data-centric tasks other than used in image classification tasks.
Although Oversampling [17] differs with the data generation, its algorithms can be employe in this area. Wei et al. [18] introduce the Improved and Random Synthetic Minority Oversampling Technique (IR-SMOTE). In data pre-processing, the K-means algorithm is utilized to cluster both majority and minority class samples, enabling the removal of noise samples in the minority class based on a distance metric. Subsequently, the kernel density estimation technique is employed to dynamically assign the number of synthetic samples to each cluster within the minority class. Finally, a synthetic approach that considers the attributes of the data is introduced to generate a more diverse set of synthetic samples.
@@ III Methodology
As shown in fig 1, the general methodology of AGCL includes feature extraction, K-means clustering for grouping data samples, and assessment of the obtained clusters to identify class representation. The cluster evaluation is based on the separation criterion, with a threshold guiding the decision to further subdivide for improved class separation. These steps pave the way for the subsequent steps in the algorithm involved in data synthesis. The analysis of each part of the methodology is presented in the subsequent sections of the paper.
@@ III-A Feature Extraction and Clustering
In the first step, relevant features are extracted from the dataset. Let $X=\{x_{1},x_{2},...,x_{n}\}$ be the dataset in the latent space. By K-means clustering, data points are organized into clusters $C=\{c_{1},c_{2},...,c_{k}\}$. The K-means algorithm operates through an iterative process aimed at minimizing the total sum of distances between each point in a cluster and its centroid [19]. By computing the mean of data points, new cluster centroids are determined, resulting in an arrangement where data points align effectively with their respective clusters. Each cluster $c_{j}$ comprises data points $x_{i,j}$ where $i=1,2,...,n_{j}$, with $n_{j}$ representing the number of data points in cluster $c_{j}$.
@@ III-B Cluster Evaluation and Re-Clustering Process
III-C Synthetic Data Generation
@@
:$:\left\{r_{j}\leftarrow\begin{cases}0.01\times\frac{1}{d}\sum_{i=1}^{d}\sigma_{j}&\text{if }n_{c_{j}}==1\\ \max(D_{x,\mu_{j}})+0.1\times(\max(D_{x,\mu_{j}})-\min(D_{x,\mu_{j}}))&\text{if }n_{c_{j}}>1\end{cases}\\ \right.$
@@
III-D Dataset Update and Classification
The dataset is updated with the synthetic data points to form $X_{augmented}\leftarrow X\cup\{x_{synth}\}$, providing an enriched set of training data for machine learning models. Leveraging machine learning models on the updated dataset enables the classification tasks for identifying patterns and making predictions.
In summary, AGCL tackles the challenges of small medical datasets by combining feature extraction, iterative K-means clustering, and synthetic data generation. The augmented dataset, comprising both original and synthetic data points, is utilized to train machine learning models, enhancing their performance in classification tasks. Ultimately, the algorithm is presented generically in Algorithm 1.
Input : Dataset $X$, Labels $y$, Parameters $\textit{thrsh},\alpha$
Output : Augmented dataset $X_{augmented}$
$X\leftarrow\text{FeatureExtractor}(X)$
Fit K-means on $X$ to obtain Clusters $C=\{c_{1},c_{2},...,c_{k}\}$ with centers $\{\mu_{1},\mu_{2},...,\mu_{k}\}$
$X_{synth}\leftarrow\emptyset$
for $j=1$ to $k$ do
       $X_{j}\leftarrow\{x\mid x\in X\text{ and }x\in c_{j}\}$
       $y_{j}\leftarrow\{y\mid x\in c_{j}\}$
       $n_{c_{j}}\leftarrow\text{size}(X_{j})$
      if $\text{Unique}(y_{j})$ is False  then
             Separation $\textit{Separation Criterion}\leftarrow\frac{\text{CSM}(c_{j})}{\text{H}(c_{j})}$
[…]
                   Call the algorithm recursively with $X_{j}$, $y_{j}$
      else
$r_{j}\leftarrow\begin{cases}0.01\times\frac{1}{d}\sum_{i=1}^{d}\sigma_{j}&\text{if }n_{c_{j}}=1\\ \max(D_{x,\mu_{j}})+0.1\times(\max(D_{x,\mu_{j}})-\min(D_{x,\mu_{j}}))&\text{if }n_{c_{j}}>1\end{cases}$
             $q\leftarrow\alpha\cdot n_{c_{j}}$
             $\Sigma_{j}\leftarrow\text{diag}(r_{j}^{2})$
            for $i=1$ to $q$ do
                   $x_{new}\sim N(\mu_{j},\Sigma_{j})$
                   $X_{synth}\leftarrow X_{synth}\cup\{x_{new}\}$
$X_{augmented}\leftarrow X\cup X_{synth}$
Algorithm 1 AGCL Pseudocode
@@ IV Dataset
Our dataset consists of videos from 55 individuals captured using an Android tablet camera. Prior to the task, participants viewed a brief video of a person performing the task. They were then asked to make a facial expression and hold it for a few seconds. The facial expressions include anger, happiness, disgust, fear, and surprise. Examples of these facial expressions are illustrated in Figure 2. Not all participants completed all five facial expressions; therefore, the distribution of videos for each emotion is presented in Table I. Since each video represents a transition from a neutral face to an extreme emotion, the peak frame of each video is selected for analysis.
Figure 2: Facial expressions from left to right: anger, happy, disgust, fear and surprise
TABLE I: Dataset Characteristics
Emotion Rage Happy Disgust Fear Surprise
Parkinson’s (PD) 26 25 26 25 29
Control 28 30 27 30 26
TABLE II: Associated AUs to each facial expression
@@
V Experiments
In this section, we evaluate the AGCL method in various scenarios. Our objective is to provide a thorough assessment of AGCL’s performance under diverse conditions.
V-A Feature Extraction and Latent Space
@@
A facial expression can be associated with multiple action units. In this paper, each facial expression is associated with three Aus by analyzing Aus’ mean, concept of each AU associated with facial expression and Pearson Correlation. Table II shows the AUs associated with facial expressions.
TABLE III: Comparison between features of real and synthetic data based on t-test, Levene-test and ks-test
t-test (p-value) Levene-test (p-value) ks-test (p-value)
@@
V-B Statistical evaluations
Three different tests were conducted to verify that the real data and synthetic data do not differ statistically. Twenty random samples were selected from each class of the two datasets (real and synthetic). The p-values for each test are shown in the table III. In the t-test [28], it is observed that the p-value is above 0.05 for all variables except one, indicating that the means of the two datasets are not significantly different. Levene’s test [29] was used to compare the variances between the datasets. The p-value is above 0.05 for all variables except one, suggesting that the variances are not statistically significant. Similarly, the most p-values from the Kolmogorov-Smirnov test (ks-test) [30] are also above 0.05, showing that the distribution of variables in the two datasets is not statistically significant. Therefore, we can conclude that the two datasets are statistically similar.
V-C Classification
TABLE IV: Comparison between accuracy of AGCL, GN, kNNMTD and Baseline on 5 different emotions.

### v2

# Feature-to-Image Data Augmentation: Improving Model Feature Extraction with Cluster-Guided Synthetic Samples

@@ Abstract
One of the growing trends in machine learning is the use of data generation techniques, since the performance of machine learning models is dependent on the quantity of the training dataset. However, in many real-world applications, particularly in medical and low-resource domains, collecting large datasets is challenging due to resource constraints, which leads to overfitting and poor generalization. This study introduces FICAug, a novel feature-to-image data augmentation framework designed to improve model generalization under limited data conditions by generating structured synthetic samples.
FICAug first operates in the feature space, where original data are clustered using the k-means algorithm. Within pure-label clusters, synthetic data are generated through Gaussian sampling to increase diversity while maintaining label consistency. These synthetic features are then projected back into the image domain using a generative neural network, and a convolutional neural network is trained on the reconstructed images to learn enhanced representations.
Experimental results demonstrate that FICAug significantly improves classification accuracy. In feature space, it achieved a cross-validation accuracy of 84.09%, while training a ResNet-18 model on the reconstructed images further boosted performance to 88.63%, illustrating the effectiveness of the proposed framework in extracting new and task-relevant features.
@@
Data augmentation, Small datasets, Parkinson’s Disease Screening, Self-Supervised Learning, Ganimation
@@ I Introduction
One of the significant challenges in machine learning is collecting high-quality, diverse, and labeled data. The performance of most models relies directly on the quality and quantity of training data [1].
In many real-world applications—especially in sensitive domains like healthcare—data are scarce, difficult to acquire, and prone to have legal, ethical, and budgetary constraints. The challenge becomes even more severe when dealing with rare diseases or specific populations, few labeled examples are available. These limitations highlight the importance of data-efficient methods that can handle well even when only a little training data is available [2, 3, 4].
To address data scarcity, various augmentation and synthetic data generation techniques have been proposed. Among them, generating new data in the feature space is a promising strategy [5]. In feature space, each data sample is mapped to a set of descriptive features, and new data points can be synthetically created by leveraging distributions that capture intra-class variability. Feature space generation provides a controlled and label-consistent sampling, making this methodology very useful for enlarging small training sets without disrupting the underlying structure of the data.
In this study, we introduce the FICAug (feature-to-image clustered augmentation) pipeline designed to improve classification performance on small datasets, with a focus on Parkinson’s disease screening. The FICAug pipneline begins by clustering the original feature space using the k-means algorithm, followed by evaluation and refinement to ensure cluster purity. Within each clusters that comes as single label, synthetic feature vectors are generated using gaussian sampling. The synthetic features are mapped from the feature-space to image domain through a generative neural network in order to push the model learn new and more discriminative features that are not present in our original feature representation.
This pipeline enables training a convolutional neural network (CNN) on the reconstructed images, leading to improved performance and generalization. By using this approach, the model learns new and informative representations from the structured synthetic data derived from existing features.
Our main objective is to demonstrate that sample from latent meaningful clusters and learning with the reconstruction from other data will enhance the performance of models. Experiments show that the proposed method improves classification performance and offers a novel solution to the challenge of learning from limited data. This paper is focused on Parkinson’s disease screening but the proposed approach is general enough to be applied in different domains with small data.
The rest of this article is organized as follows. In Section II, we provide an overview of existing research on data augmentation. Section III details our proposed method, including the clustering-based feature augmentation process and the generative reconstruction strategy used to produce image-space data. Section IV introduces the dataset used in our experiments, along with its characteristics and preprocessing steps. Section V presents the experimental setup and performance analysis. Finally, Section VI summarizes the key findings.
@@ II Related Work
Small dataset learning is different from the closely related task of transfer learning, few-shot learning and semi-supervised methods. Unlike transfer learning, where models benefit from large-scale pretraining, or few-shot learning, where auxiliary classes with many labeled samples support generalization, small-data learning typically lacks access to both peripheral data and domain-specific priors. Semi-supervised learning, on the other hand, assumes the availability of a large pool of unlabeled data—which is often not the case in sensitive applications like medicine [6, 7, 8].
@@
When it comes to working with small datasets, data augmentation appears to be one of the typical methods to improve the performance of machine learning models. There are many augmentation methods used for image classification, including random rotation, mirroring, and the addition of Gaussian noise [9]. Heuristic methods such as Cutout [10] partially occlude a square area in each training process to affect the learned features, and Random Erasing [11] partially covers or replaces certain areas of an image.
Mixing is another significant method of data augmentation that involves combining images or regions of interest within a single image. These methods typically operate by combining two or more images—or subregions of them—into a single composite image. One widely used approach is Mixup [12], which generates weighted combinations of random image pairs by linearly interpolating both their pixel values and labels. Another method, known as CutMix [13], replaces a randomly selected region of an image with a patch from another image.
Recent advancements in automatic augmentation aim to optimize augmentation strategies using reinforcement learning, offering potential efficiency improvements [14]. However, these methods often face high computational costs due to the complexity of searching for optimal policies. Alternatively, feature augmentation techniques such as FitMatch [15] generate new representations directly in feature space. While effective, they may significantly increase data dimensionality and model complexity, leading to longer training times.
Generative adversarial networks (GANs) offer a promising solution for augmenting datasets, especially with the growing availability of powerful generative models [16]. However, their effectiveness is often limited in low-data settings due to the large sample size typically required for training. To address this, the Least-Square GAN (LS-GAN) was introduced. It starts with feature decorrelation using statistical tests, followed by Burato’s feature selection method based on random forests. LS-GAN also employs a least-squares loss to prevent gradient vanishing and improve training stability [17].
Feature space data augmentation has become an effective strategy for small datasets because it offers more control over data distribution than raw input-space augmentation. A simple approach in this domain is perturbation, where techniques like Gaussian noise injection simulate variability in a controlled way [18]. For more structure, the kNNMTD method [19] generates synthetic data using $k$-nearest neighbors and Mega-Trend Diffusion. It selects neighbor points and uses membership functions to ensure synthetic samples preserve the original data relationships.
Moreover, MEML [20] addresses deep learning contexts by generating synthetic features within intermediate network layers. It computes class-wise feature means and extrapolates between them and actual samples to produce meaningful representations during training. Autoencoders are also used for synthetic data generation. DeVries and Taylor [21] apply transformations like Gaussian noise, interpolation, and extrapolation to learned context vectors. This allows the creation of diverse synthetic samples while maintaining semantic alignment with the original data.
Self-supervised learning methods have emerged as powerful tools for extracting rich representations without relying on labeled data. Two prominent examples are DINO [22] and Masked siamese network(MSN) [23]. DINO leverages Vision Transformers in a self-distillation framework, encouraging different views of the same image to produce similar embeddings through teacher–student networks. This allows the model to capture diverse and meaningful features applicable to downstream tasks without supervision.
Similarly, MSN proposes a Masked Siamese Network that learns representations by comparing masked views of images through a contrastive-style architecture, emphasizing label-efficient learning. Both DINO and MSN are self-supervised learning frameworks designed to extract general-purpose features from large-scale unlabeled datasets. These features are often used in downstream tasks with or without fine-tuning.
@@ III Methodology
Figure 1: Overview of the FICAug framework, which generates synthetic feature vectors through clustering and Gaussian sampling, reconstructs them into images via a generative model, and trains a CNN to extract enhanced task-specific features.
As shown in Figure 1, the general methodology of FICAug includes generating synthetic data points in the feature space by applying k-means clustering to the original dataset and identifying label-consistent regions. Within these clusters, new samples are created using Gaussian sampling to enhance intra-class diversity while maintaining the underlying distribution. These synthetic features are then projected back into the image domain using a generative neural network. A convolutional neural network (CNN) is trained on these reconstructed images, enabling the model to extract richer and more task-specific features from structured synthetic data.
@@ III-A Feature Extraction and Clustering
In the first step, the relevant features are extracted from the dataset. Let $X=\{x_{1},x_{2},...,x_{n}\}$ be the dataset in the latent space. By K-means clustering, data points are organized into clusters $C=\{c_{1},c_{2},...,c_{k}\}$. The K-means algorithm operates through an iterative process aimed at minimizing the total sum of distances between each point in a cluster and its centroid [24]. By computing the mean of data points, new cluster centroids are determined, resulting in an arrangement where data points align effectively with their respective clusters. Each cluster $c_{j}$ comprises data points $x_{i,j}$ where $i=1,2,...,n_{j}$, with $n_{j}$ representing the number of data points in cluster $c_{j}$.
@@ III-B Cluster Evaluation and Re-Clustering Process
III-C Synthetic Data Generation in Feature Space
@@
$r_{j}\leftarrow\begin{cases}0.01\times\frac{1}{d}\sum_{i=1}^{d}\sigma_{j}&\text{if }n_{c_{j}}==1\\ \max(D_{x,\mu_{j}})+0.1\times(\max(D_{x,\mu_{j}})-\min(D_{x,\mu_{j}}))&\text{if }n_{c_{j}}>1\end{cases}\$
@@
III-D Mapping Synthetic Data to Original Space
The synthetic data generated in the feature space are mapped back to the original space using a reconstruction network. This network is designed to interpret the abstract features of the synthetic data and transform them into a format that closely resembles the original dataset.
The reconstruction process ensures that the synthetic data retain the structural and contextual characteristics of the real data, making it suitable for further analysis and training. By mapping the synthetic data to the original space, the model can leverage the augmented dataset to enhance its learning capabilities while maintaining the real-world properties of the data.
III-E CNN Model Training and Fine-Tuning
The Convolutional Neural Network (CNN) is initially trained on the synthetic dataset, which includes only the reconstructed synthetic data. This training phase allows the model to learn robust features from the enriched dataset, improving its ability to generalize across diverse data samples.
Once the initial training is complete, the CNN is fine-tuned on the original dataset to adapt to the specific nuances and patterns present in the real data. Fine-tuning ensures that the model retains its generalization capabilities while optimizing its performance for the target task.
Finally, the trained and fine-tuned CNN performs feature extraction and final classification, using the learned features to accurately identify patterns and make predictions. This integrated approach enhances the model’s performance, particularly in scenarios with limited data, by combining the benefits of synthetic data augmentation and targeted fine-tuning.
@@ IV Dataset
Our dataset consists of videos from 55 individuals captured using an Android tablet camera. Prior to the task, participants viewed a brief video of a person performing the task. They were then asked to make a facial expression and hold it for a few seconds. The facial expressions include anger, happiness, disgust, fear, and surprise. Examples of these facial expressions are illustrated in Figure 2. Since each video represents a transition from a neutral face to an extreme emotion, the peak frame of each video is selected for analysis.
Figure 2: Facial expressions from left to right: anger, happiness, disgust, fear and surprise
V Experiments
In this section, we evaluate the effectiveness of the proposed FICAug framework across multiple experimental settings. We compare FICAug with baseline models and existing augmentation strategies in both feature space and image space, highlighting its capability to generate informative synthetic samples and improve feature learning.
V-A Feature Extraction and Feature Space
TABLE I: Associated AUs to each facial expression
@@ V-A Feature Extraction and Feature Space
A facial expression can be associated with multiple action units. In this paper, each facial expression is associated with three Aus by analyzing Aus’ mean, concept of each AU associated with facial expression and Pearson Correlation. Table I shows the AUs associated with facial expressions.
V-B Data Splitting for Training, Validation, and Testing
We employed a Two-Leave-Out cross-validation approach, creating 25 unique folds where each iteration used one control and one patient sample for validation while training on the remainder. This ensures comprehensive, unbiased evaluation as all samples serve in validation exactly once. The data partitioning maintains balanced age and gender distributions across sets to minimize bias. Final predictions combine results from all 25 models through weighted majority voting, where each model’s contribution reflects its validation performance.
TABLE II: Distribution of training, validation and test sets for each emotion
Dataset Group Anger Happiness Disgust Fear Surprise
Train/Val Parkinson 22 18 23 22 22
Control 19 19 17 19 17
Test Parkinson 4 7 3 2 7
Control 9 11 11 11 9
V-C Statistical evaluations
TABLE III: Statistical comparison between features of real and synthetic data using t-test, Levene’s test, and Kolmogorov–Smirnov (KS) test, with corresponding p-values.
t-test Levene-test ks-test
@@ V-C Statistical evaluations
Three different tests were conducted to verify that the real data and synthetic data in feature space do not differ statistically. Twenty random samples were selected from each class of the two datasets (real and synthetic). The p-values for each test are shown in the table III. In the t-test [33], it is observed that the p-value is above 0.05 for all variables except one, indicating that the means of the two datasets are not significantly different. Levene’s test [34] was used to compare the variances between the datasets. The p-value is above 0.05 for all variables except one, suggesting that the variances are not statistically significant. Similarly, the most p-values from the Kolmogorov-Smirnov test (ks-test) [35] are also above 0.05, showing that the distribution of variables in the two datasets is not statistically significant. Therefore, we can conclude that the two datasets are statistically similar.
V-D Evaluation of Feature Space Augmentation Strategies
The goal of this experiment is to evaluate whether the proposed augmentation strategy in the feature space is effective enough to serve as a reliable foundation for our next steps, including image generation and training CNNs for feature learning. We aim to determine whether FICAug can produce meaningful and diverse synthetic samples that perform competitively in classification tasks, making it a valid alternative to other augmentation methods.
To this end, we compared FICAug against three approaches: a baseline (no augmentation), Gaussian noise (GN) perturbation, and the $k$NNMTD algorithm [19], using four classifiers across five emotions. Gaussian noise augmentation involved adding random noise drawn from a normal distribution with varying standard deviations. For FICAug, synthetic data points were generated within well-separated clusters using a normal distribution guided by cluster-specific parameters, with grid search used to tune the $\alpha$ value controlling augmentation size.
As shown in Table IV, the results across validation and test accuracy demonstrate that FICAug generally performs on par with GN and $k$NNMTD, occasionally outperforming them in specific emotions. In Rage, FICAug reached a notable test accuracy of 82.82% with MLP. The relatively high standard deviation in the validation results is due to the two-leave-out cross-validation strategy. With only two samples (one per class) used in each validation fold, even a single misclassification can have a substantial impact on the accuracy, resulting in greater variability across folds.
For Happiness and Disgust, proposed method maintained strong and competitive performance. In Happiness, it reached a high validation accuracy of 81.25% with both MLP and Random Forest, and a top test accuracy of 83.33% using SVM. In Disgust, the FICAug again performed best in test accuracy (64.28%) despite Gaussian noise occasionally showing slightly better validation results. These findings indicate that while some methods show short-term gains, the proposed strategy generalizes better to unseen data.
In the Fear emotion, although our method achieved the highest validation accuracy (76.00%) with MLP, its final test results were comparable to those of Gaussian noise. In Surprise, FICAug achieved 75.00% in test and 82.00% in cross validation, demonstrating its effectiveness in generating high-quality augmented data. However, when aggregating predictions across all emotions using majority voting, the FICAug delivered the most consistent and accurate overall results. It achieved the highest test accuracy across all classifiers, and attained the top validation score of 84.09% with Random Forest. The final test accuracy reached 77.77% (RF) and 83.33% (MLP), showing significant improvement over the baseline.
While FICAug does not always significantly outperform other methods in classification accuracy in feature space, its purpose is not solely to boost performance at the feature level. Instead, the proposed augmentation strategy in feature space is designed as a structured augmentation mechanism that avoids generating synthetic points in mixed or ambiguous regions of the feature space—areas where different classes intersect. This selective sampling is crucial for guiding the model to focus on distinct, class-specific regions, enabling more effective downstream learning. Unlike generic noise injection or diffusion-based sampling, FICAug incorporates a class-separation-aware strategy, which we believe is key to generating training data that leads to more meaningful representations when mapped back to the image space.
TABLE IV: Comparison between accuracy of FICAug, GN, kNNMTD and Baseline on 5 different emotions.